Two train are 532 miles apart and are traveling toward each other. One train is traveling 9 mph faster than the other train. The trains pass each other in 4 hours. Fine the speed of each train.
Answer by mananth(16946)

Two train are 532 miles apart and are traveling toward each other. One train is traveling 9 mph faster than the other train. The trains pass each other in 4 hours. Fine the speed of each train.
..
let one train's speed be x
other speed = x+9
..
since they are travelling towards each other the effective speed =2x+9 mph
distance = 532 miles
time = distance/speed
532/(2x+9)=4
532=4*(2x+9)
532=8x+36
8x=532-36
8x=496
/8
x=62 mph speed of one train
other train speed = 62+9 = 71 mph
...
m.ananth@hotmail.ca